# EpidemIQs: Prompt-to-Paper LLM Agents for Epidemic Modeling and Analysis

### v1

@@ V-C Discussion, Limitations, and Ethical Concerns
Our framework is unique, with no direct comparators. However, a comparison with The AI Scientist, a general tool for scientific discovery from ideation to paper writing [34], for question one, showed the failure of The AI Scientist. It analyzed the ODE-based analytical component but struggled with stochastic simulations over the network, producing hallucinated results and conclusions. It is essential to mention The AI Scientist performance was mostly evaluated on AI-related subjects, such as diffusion models or transformers, emphasizing that this comparison does not imply overall superiority of one framework over another.
Summarizing the main limitations of the current work, we have:
@@
  • Although no major hallucinations were found in the results, a few hallucinated references did appear in the final report. Moreover, while the reviewers praised the papers’ comprehensiveness, they also noted some repetitive content that lengthened the reading time, and sometimes tables were so large that portions extended beyond the page, suggesting enhancing the report generation part.
@@
  • The performance of the EpidemIQs is highly dependent on the LLM models, while our framework is model agnostic, as long as these models can support structured outputs and tool calling, the results can vary based on the LLM choices.
@@
  • The literature review relies only on the abstract and summary of key findings, rather than delving deeply into the full content of the papers. Future frameworks capable of thoroughly analyzing the full material of each paper could enhance and benefit from deeper data retrieval.
  • The current framework focuses on epidemic modeling over complex networks as a proof of concept. Enhancing it to broader methods such as Agent-Based Models(ABM)[60, 61], individual-based [60], or statistical and data-driven approaches [60, 62] can significantly increase the applicability of the framework.
@@
  • The agents’ performance hinges on the tools at their disposal. As noted in question three, their performance and reliability decline when they encounter highly technical problems without the requisite expert knowledge and appropriate tools. This shortcoming explains why the general‑purpose module proposed by [34] fails in question one.
@@
  • AI evaluation in the automated review process shows divergence from human expert reviews, aligning with other works’ findings, such as Agent Laboratory [30]. However, contrasting with near-human performance reported in [34], it suggests that human involvement is necessary. While an LLM-as-Judge can be beneficial for providing high-level insights, it cannot be fully trusted, since it relies on superficial patterns rather than robust analysis criteria.
@@
  • We did not include ideation in our framework, primarily because epidemic research is highly complex and interdisciplinary. Generating new research ideas requires careful consideration of resources, advanced tools (see Limitation 4), expert knowledge across multiple domains, and access to diverse data sources, such as mobility, weather, and epidemiological data. For example, in vector-borne diseases, the populations of mosquitoes or birds play a critical role in epidemiological modeling. To the authors’ knowledge, such ideas can quickly become too complex for current agentic AI systems to handle appropriately, given concrete bottlenecks such as the lack of structured access to high-quality mobility/weather/entomological data, difficulty in reasoning across coupled mechanistic models, or access to specific tools. Therefore, we focused on specific network-based epidemic modeling tasks as a proof of concept to showcase the framework’s potential benefits.
Therefore, although EpidemIQs offers great potential to accelerate epidemic-modeling research and allows researchers to focus on high-level ideation, this promise also introduces new ethical challenges that must be taken seriously. Its ability to autonomously generate high-quality papers, supported by analytical insights, rigorous simulations, and comprehensive analyses, may inadvertently lower the barriers to producing low-standard scientific work, potentially overwhelming the peer-review system and jeopardizing the integrity of academic discourse. Moreover, as noted earlier, the automated process may introduce mistakes in algorithm design that are implicit and difficult to detect, or it may yield results biased by its underlying knowledge or algorithms, leading to skewed outcomes. These concerns underscore the necessity of disclosing AI involvement in the research process to mitigate such risks and maintain accountability, fairness, and transparency.
@@ VI Data Availability
All the data and results discussed in this paper are provided as supplementary materials. The EpidemIQs is also provided as an open-source Python module at github.com/will-be-announced-upon-publication.
@@ VII Conclusion
In this study, we present EpidemIQs, a multi‑LLM agent framework that harnesses the reasoning, planning, and problem‑solving abilities of LLMs and integrates them with online and literature retrieval, multimodal data processing, code generation and execution, stochastic simulation, and visual and data analysis. These characteristics enable it to autonomously carry out every key stage of epidemic modeling research, from discovery through modeling and simulation to analysis and, finally, the synthesis of findings into a full scientific report. This entire process takes less than half an hour, whereas it would normally take an average graduate-level student days to accomplish the same task.
@@
The comparative analysis of EpidemIQs and the single-agent implementation consistently showed that EpidemIQs outperformed the single-agent variants across all tasks. It achieved an 100% success rate average human‑evaluation score of 7.98 at a cost of \$1.57, whereas the single-agent configurations based on GPT‑4.1 and o3 obtained scores of 5.06 and 5.68 at costs of \$0.91 and \$4.13, respectively.
At present, EpidemIQs focuses exclusively on epidemic modeling over networks as a proof of concept; nevertheless, it has already demonstrated the potential of deploying LLM agents in a collaborative environment. This opens an exciting path for further performance gains by integrating advanced tools, such as deep learning techniques, probabilistic methods, and data-centric models, heralding a promising future for epidemic modeling research. Moreover, because the framework’s accuracy and performance depend heavily on the underlying LLMs, the rapid advances in this field are likely to yield substantial improvements for the entire multi agent system, as our framework is designed to be model agnostic.
@@
Future work should consider the broader applications of epidemic modeling by integrating real-world data and other advanced computational tools into the model design to address open-ended real-world problems and evaluate its applicability for forecasting current outbreaks. Additionally, they should address ethical concerns regarding the misuse or dual use of such a framework, as automated modeling and report generation could be exploited to produce misleading forecasts or scientifically plausible misinformation, intentionally or unintentionally influencing public perception or policy decisions. The system’s ability to simulate outbreak dynamics at scale may also lower barriers for malicious actors to explore harmful scenarios or identify vulnerabilities in public health defenses. Additionally, integration with sensitive epidemiological data poses privacy risks, emphasizing the need for implementing appropriate safeguards.

### v2

@@ V-D Limitations and Future Directions
To the best of the authors’ knowledge, our framework is unique, with no direct comparators. However, a comparison in Question 1 with The AI Scientist, a general tool for scientific discovery from ideation to paper writing [31], showed that it was unable to successfully address the task. It analyzed the ODE-based analytical component but struggled with stochastic simulations over the network, producing hallucinated results and conclusions. The Jupyter Agent 2, as an AI coding agent, fell short in addressing analytical aspects. In contrast, Denario [48] showed better performance on analytical tasks but struggled with stochastic simulations and threshold vaccination in Question 5. It is essential to note that the performance of these AI tools was primarily evaluated in areas other than epidemics, including diffusion models and transformers, coding tasks, and biomedical tasks. This emphasizes that this comparison does not imply overall superiority of one framework over another, but highlights the unique benefits our framework can offer. Summarizing the main limitations of the current work, we have:
@@
  • Although no major hallucinations were found in the results, a few hallucinated references did appear in the final report. Moreover, while the reviewers praised the papers’ comprehensiveness, they also noted repetitive content that lengthened reading time and, at times, tables that were large, with portions extending beyond the page, suggesting a need to improve the report-generation process.
@@
  • The performance of the EpidemIQs is highly dependent on the LLM models. While our framework is model-agnostic, as long as these models can support structured outputs and tool calling, the results will vary depending on the LLM used. Since LLMs are word sensitive, a promising future direction is to use the current multi-agent framework as a teacher model for prompt optimization when integrating new LLMs.
@@
  • The literature review relies only on the abstract and summary of key findings, rather than delving deeply into the full content of the papers. Future frameworks capable of thoroughly analyzing the full content of each paper could benefit from deeper data retrieval.
  • The current framework focuses on epidemic modeling over complex networks as a proof of concept. Extending it to broader methods, such as Agent-Based Models (ABMs), individual-based models, or statistical and data-driven approaches [56, 57], can significantly increase the applicability of the framework. Such extensions can be enabled by equipping the modeling and simulation phases with new tool interfaces and corresponding one-shot examples, allowing the framework to replace FastGEMF with alternative engines. In addition, running multiple modeling pipelines in parallel, where each is based on a different modeling paradigm, would produce an ensemble of outcomes, potentially improving robustness and reducing uncertainty in the final results.
@@
  • The agents’ performance hinges on the tools at their disposal. As noted in Question 3, with the lowest task success rate of 65%, their performance and reliability decline when they encounter highly technical problems without the requisite expert knowledge and appropriate tools. This shortcoming explains why the general‑purpose module proposed by [31] fails in Question 1.
@@
  • AI evaluation in the automated review process shows divergence from human expert reviews, aligning with other works’ findings, such as Agent Laboratory [27]. However, in contrast to the near-human performance reported in [31], this suggests that human involvement is necessary. While an LLM-as-Judge can be beneficial for providing high-level insights, it cannot be fully trusted, since it relies on superficial patterns rather than robust analysis criteria. Furthermore, AI evaluation shows a lower tendency toward harsh penalization of incorrect logic or weak experimental rigor than human experts, who usually place greater weight on internal consistency and methodological soundness. Another reason is that AI assessments rely solely on the manuscript text, whereas human reviewers often verify the reported results against the generated scripts and outputs, allowing them to assess whether the findings align with the claimed results.
  • While our multi-faceted evaluation system (Section IV) provides a practical approach that could inform future standardization efforts, there is a lack of widely accepted, statistically grounded benchmarks for end-to-end agentic systems, specifically in the epidemic field. Addressing shared tasks, scoring rubrics, and statistically robust evaluation protocols is an important limitation and a direction for future work.
@@
  • We did not include ideation in our framework, primarily because epidemic research is highly complex and interdisciplinary. Generating new research ideas requires careful consideration of resources, advanced tools (see Limitation 4), expert knowledge across multiple domains, and access to diverse data sources, such as mobility, weather, and epidemiological data. For example, in vector-borne diseases, mosquito or bird populations play a critical role in epidemiological modeling. To the authors’ knowledge, such ideas can quickly become too complex for current agentic AI systems to handle appropriately, given concrete bottlenecks such as the lack of structured access to high-quality mobility/weather/entomological data, difficulty in reasoning across coupled mechanistic models, or access to specific tools. Therefore, we focused on specific network-based epidemic modeling tasks as a proof of concept to showcase the framework’s potential benefits.
V-E Ethical Concerns and Challenges
Therefore, although EpidemIQs demonstrated the great potential of agentic AI to accelerate epidemic modeling research and allow researchers to focus on high-level ideation, this promise also introduces new ethical challenges that must be taken seriously. Specifically:
  • Scientific Misinformation: With easy access to advanced modeling tools, the primary risk is the production of incorrect or low-quality papers. As our evaluation demonstrated, EpidemIQs, like any other AI system [49, 48, 34], can produce outputs that appear plausible and methodologically sound in structure and tone but contain errors (Section V-D). For example, as detailed in our failure analysis (Section V-C1), an algorithmic flaw in one trial led to the premature recovery of infected nodes, resulting in a significant underestimation of the final outbreak size. In a real-world policy setting, overreliance on such a plausible-sounding but technically flawed conclusion could lead to the premature relaxation of public health interventions, thereby increasing overall risk to the population. The other example was when the iterative single-agent incorrectly concluded that herd immunity through high-degree vaccination cannot suppress the disease because of an incorrect assumption about the contact-degree distribution. This highlights the risk of misleading persuasiveness, where the AI’s confident tone in its conclusion masks critical underlying mathematical errors that can lead to public misinformation.
    Furthermore, the low computational cost of generating a paper for ($\approx\mathdollar 1.57$) poses a risk of flooding the scientific record with low-quality or fabricated studies. If unchecked, this could overwhelm the peer-review system, creating a denial-of-service effect on human experts who must verify these outputs. While efforts have been made to detect AI-generated text with high confidence, AI detection remains challenging due to the complexity and continuous evolution of LLMs [58].
  • Dual-Use and Biosecurity Risks: In the domain of epidemiology, automated modeling tools possess inherent dual-use risks. For example, in the threshold vaccination case, the same reasoning capabilities might, in theory, be inverted by malicious actors to identify optimal conditions for pathogen spread. Although EpidemIQs is currently constrained to standard epidemic models as a proof-of-concept and does not have access to health datasets, future iterations integrated with real-world mobility or health-related data must be gated by strict access controls to prevent the automated optimization of harm. Some guardrails are already built into LLMs to prevent misuse for potentially harmful purposes. However, this trade-off limits the benefits of capable AI models for legitimate, beneficial purposes. A more sustainable solution would be selective, credentialed access to capable models for safety-critical research.
  • Authorship: The emergence of AI tools introduces new complexities to scientific authorship. While AI may generate content, accountability for scientific validity can not be delegated to the agentic system. Authorship must be shifted from content generation to rigorous verification. EpidemIQs is designed strictly as an AI research assistant subject to human oversight, assuming no authorship rights. Human researchers must explicitly state the use of agentic systems and are responsible for auditing the outputs, verifying assumptions, and validating artifacts and conclusions. Therefore, responsibility and authorship remain with human researchers who must audit and verify the AI system’s outputs, including simulations, derivations, and logical reasoning of the AI framework.
To mitigate some of these risks, we have implemented automatic watermarking to prevent authorship confusion in every generated report and a restrictive usage policy that prohibits removing watermarks or submitting EpidemIQs outputs to scientific venues without significant human contribution and verification. Users must explicitly acknowledge that the system is a research copilot subject to execution errors. A SecretaryAgent is implemented as a classifier to trigger EpidemIQs only if the query is safe and within the scope of the framework as an effort to reduce the chance of malicious activity or hallucination.
@@ VI Data Availability
All the data and results discussed in this paper are provided in Supplementary Sections LABEL:S-sec:q1 to LABEL:S-sec:q10 along with reports in Section LABEL:S-sec:papers, and the project repository. EpidemIQs is provided as an open-source Python module at github.com/KsuNetse/EpidemIQs. All generated results, configurations of EpidemIQs for five epidemic scenarios discussed here, and five additional cases are available at this repository.
@@ VII Conclusion
In this study, we present EpidemIQs, a multi‑LLM agent framework that harnesses the reasoning, planning, and problem‑solving abilities of LLMs and integrates them with online and literature retrieval, multimodal data processing, code generation and execution, stochastic simulation, and visual and data analysis. These characteristics, along with its proposed orchestration, enable it to autonomously execute every key stage of the pipeline, from discovery through modeling and simulation to analysis, and finally synthesize the findings into a scientifically structured report summarizing the workflow outcomes. This entire process takes less than half an hour, whereas a human expert would require at least 38.4 hours to complete an execution cycle.
@@
The comparative analysis of EpidemIQs and the single-agent implementation consistently showed that EpidemIQs outperformed the single-agent variants across all tasks. It achieved a 100% workflow completion rate with an average human‑evaluation score of 7.98 at a cost of \$1.57, whereas the single-agent configurations based on GPT‑4.1 and o3 obtained scores of 5.06 and 5.68 at costs of \$0.91 and \$4.13, respectively.
At present, EpidemIQs focuses exclusively on epidemic modeling over networks as a proof of concept; nevertheless, it has already demonstrated the potential of deploying LLM agents in a collaborative environment. This opens an exciting path for further performance gains by integrating advanced tools, such as deep learning techniques, probabilistic methods, and data-centric models, heralding a promising future for epidemic modeling research. Moreover, because the framework’s accuracy and performance depend heavily on the underlying LLMs, the rapid advances in this field are likely to yield substantial improvements for the entire multi-agent system, as our framework is designed to be model agnostic.
@@
Future works should consider the broader applications of epidemic modeling by integrating real-world data and other advanced computational tools into the model design to address open-ended real-world problems and evaluate its applicability for forecasting current outbreaks. Additionally, future work should address ethical concerns regarding the misuse or dual use of such a framework, as automated modeling and report generation could be exploited to produce misleading forecasts or scientifically plausible misinformation, intentionally or unintentionally influencing public perception or policy decisions. The system’s ability to simulate outbreak dynamics at scale may also lower barriers for malicious actors to explore harmful scenarios or identify vulnerabilities in public health defenses. Additionally, integration with sensitive epidemiological data poses privacy risks, emphasizing the need for implementing appropriate safeguards.